Find the class boundaries midpoint and width for the class 47-55


1
Expert's answer
2021-06-18T10:56:41-0400

The size of the gap between classes is the difference between the upper class limit of one class and the lower class limit of the next class. In our case, gap "=1."

The lower boundary of each class is calculated by subtracting half of the gap value from the class lower limit. On the other hand, the upper boundary of each class is calculated by adding half of the gap value to the class upper limit.


"47-0.5(1)=46.5"

"55+0.5(1)=55.5"

The class boundaries: "46.5-55.5."

Midpoint: "\dfrac{47+55}{2}=51"

The width of class is the difference between the upper boundary and the lower boundary


"55.5-46.5=9"


The class width : "9."